jamal works part time earning 8000 per year. he reduces his hours by half so he can go to school. Jamal completes his degree after 3 semesters at the cost of $1400 per semester in tuition and $500 per semester in books. whats jamals investment in getting his degree?
1
Expert's answer
2020-09-09T17:39:03-0400

Step-by-step explanation:

Given,

His earning before course = $ 8000 per year,

According to the question,

During the course, his earning =  = $ 4000 per year,

So, the remaining earning that he lose during course = 8000 - 4000 = $ 4000,

Now, the fees of course = tuition fee of 1,400 per semester + semester fee of $ 500 per semester

If the course has 3 semester,

Then the total fees of the course = 3 × ( 1400 + 500 )

= 3 × 1900

= $ 5700,

Hence, his investment in getting the degree = 5700 + earning for 1.5 year

= 5700 + 1.5 × 4000

= 5700 + 6000

$ 11,700